# Bi-Hölder invariants in o-minimal structures

### v2

@@ 5. Homotopy types of the links of the tangent cones
By Hardt’s triviality theorem, there exists $\delta>0$ such that $g$ is a definably trivial over $(0,\delta)$. Consequently, for $0<\varepsilon_{1}<\varepsilon_{2}<\delta$, there is a strong deformation retract from $g^{-1}([0,\varepsilon_{2}]$ onto $g^{-1}([0,\varepsilon_{1}])$ (see Observation 2.9), hence the inclusion $g^{-1}([0,\varepsilon_{1}])\hookrightarrow g^{-1}([0,\varepsilon_{2}])$ induces isomorphisms on all homotopy groups.
@@
Although $\Omega_{Z}$ need not be definable in the original structure, it is definable in an exponential o-minimal expansion (see [MS2002]). In particular, $\rho_{Z}$ is definable on the same structure as $\Omega_{Z}$. Applying Hardt’s triviality theorem to $\rho_{Z}|_{\Omega_{Z}\setminus Z}$, we obtain $\delta_{Z}>1$ such that $\rho_{Z}$ is definably trivial over $(1,\delta_{Z})$. By Observation 2.9, for any $1<d<d^{\prime}<\delta_{Z}$, there exist strong deformation retract from $\rho_{Z}^{-1}([d,\delta_{Z}])=ST_{d}(Z,C)\setminus X$ onto $\rho_{Z}^{-1}([d^{\prime},\delta_{Z}])=ST_{d^{\prime}}(Z,C)\setminus Z$. Since $Z\subset ST_{d}^{\prime}(Z,C)$, we may extend this deformation by defining it to be the identity on $Z$. This yields a strong deformation retract from $ST_{d}(Z,C)$ onto $ST_{d^{\prime}}(Z,C)$. This deformation retract fixes $Z$, hence fixes 0 as well. Consequently, it induces a strong deformation retract from $ST_{d}(Z,C)\setminus\{0\}$ onto $ST_{d^{\prime}}(Z,C)\setminus\{0\}$.
@@
Since $\iota_{i,*}$, $i=1,2$ are isomorphic, so are the others in the diagram. As $\varphi$ is a homeomorphism, it follows that the homotopy groups of $ST_{d_{3}}(X,C)\setminus\{0\}$ and $ST_{d_{1}}(Y,C)\setminus\{0\}$ are isomorphic. Thus, the homotopy types of $ST_{d^{\prime}}(X,C)\setminus\{0\}$ and $ST_{d^{\prime\prime}}(Y,C)\setminus\{0\}$ are the same for all $d^{\prime},d^{\prime\prime}\in(1,d_{0})$. It is known that $ST_{d}(Z,C)\setminus\{0\}$, $Z=X,Y$ can always deformation retracts onto ${\rm Lk}(ST_{d}(Z,C),0)$. Therefore, ${\rm Lk}(ST_{d}(X,C),0)$ and ${\rm Lk}(ST_{d}(Y,C),0)$ are of the same homotopy type.
@@
By Lemma 5.1, the homotopy groups of ${\rm Lk}(ST_{d}(C_{0}(X),C),0)$ and ${\rm Lk}(C_{0}(X),0)$ are isomorphic, and likewise the homotopy groups of ${\rm Lk}(ST_{d}(C_{0}(Y),C),0)$ and ${\rm Lk}(C_{0}(Y),0)$ are isomorphic. It follows that ${\rm Lk}(C_{0}(X),0)$ and ${\rm Lk}(C_{0}(Y),0)$ have the same homotopy type.
@@ 6. Applications
In the following we present a simple proof of this theorem in a slightly stronger form.

### v3

@@ 5. Homotopy types of the links of the tangent cones
By Hardt’s triviality theorem, there exists $\delta>0$ such that $g$ is a definably trivial over $(0,\delta)$. Using Observation 2.9 (2) we see that for $0<\varepsilon_{1}<\varepsilon_{2}<\delta$, there is a strong deformation retract from $g^{-1}([0,\varepsilon_{2}]$ onto $g^{-1}([0,\varepsilon_{1}])$, hence the inclusion $g^{-1}([0,\varepsilon_{1}])\hookrightarrow g^{-1}([0,\varepsilon_{2}])$ induces isomorphisms on all homotopy groups.
@@
Although $\Omega_{Z}$ need not be definable in the original structure, it is definable in an exponential o-minimal expansion (see [MS2002]). In particular, $\rho_{Z}$ is definable on the same structure as $\Omega_{Z}$. Applying Hardt’s triviality theorem to $\rho_{Z}|_{\Omega_{Z}\setminus Z}$, we obtain $\delta_{Z}>1$ such that $\rho_{Z}$ is definably trivial over $(1,\delta_{Z})$. By Observation 2.9, for any $1<d<d^{\prime}<\delta_{Z}$, there exist strong deformation retract from $(\rho_{Z}|_{\Omega_{Z}\setminus Z})^{-1}([d,\delta_{Z}])=ST_{d}(Z,C)\setminus X$ onto $(\rho_{Z}|_{\Omega_{Z}\setminus Z})^{-1}([d^{\prime},\delta_{Z}])=ST_{d^{\prime}}(Z,C)\setminus Z$. Since $Z\subset ST_{d}^{\prime}(Z,C)$, we may extend this deformation retract by defining it to be the identity on $Z$. This yields a strong deformation retract from $ST_{d}(Z,C)$ onto $ST_{d^{\prime}}(Z,C)$ that fixes $Z$, hence fixes 0 as well. Consequently, it induces a strong deformation retract from $ST_{d}(Z,C)\setminus\{0\}$ onto $ST_{d^{\prime}}(Z,C)\setminus\{0\}$ ($\dagger$).
@@
Since $\iota_{i,*}$, $i=1,2$ are isomorphic, so are the others in the diagram. As $\varphi$ is a homeomorphism, it follows that the homotopy groups of $ST_{d_{3}}(X,C)\setminus\{0\}$ and $ST_{d_{1}}(Y,C)\setminus\{0\}$ are isomorphic. Together with ($\dagger$), we see that the homotopy types of $ST_{d^{\prime}}(X,C)\setminus\{0\}$ and $ST_{d^{\prime\prime}}(Y,C)\setminus\{0\}$ are the same for all $d^{\prime},d^{\prime\prime}\in(1,d_{0})$. It is known that $ST_{d}(Z,C)\setminus\{0\}$, $Z=X,Y$ can always deformation retracts onto ${\rm Lk}(ST_{d}(Z,C),0)$. Therefore, ${\rm Lk}(ST_{d}(X,C),0)$ and ${\rm Lk}(ST_{d}(Y,C),0)$ are of the same homotopy type.
@@
By Lemma 5.1, the homotopy groups of ${\rm Lk}(ST_{d}(C_{0}(X),C),0)$ and ${\rm Lk}(C_{0}(X),0)$ (resp. ${\rm Lk}(ST_{d}(C_{0}(Y),C),0)$ and ${\rm Lk}(C_{0}(Y),0)$) are isomorphic. It follows that ${\rm Lk}(C_{0}(X),0)$ and ${\rm Lk}(C_{0}(Y),0)$ have the same homotopy type.
@@ 6. Applications
In the following we present a simple proof of this theorem in a stronger form.